HI Mapping of Luminous IR Galaxies

J. E. Hibbard, M. S. Yun NRAO

Summary

We have mapped the HI distribution in the 7 nearest and brightest infrared (IR) galaxies using the C- and D-array configurations of the Very Large Array (VLA). All 7 galaxies (Arp 220, Mrk 273, Arp 299=NGC 3690/IC694, NGC 1614, NGC 2623, VV 114 & IC 883) are well-studied mergers of gas rich systems. We compare the HI content, distribution and kinematics of these IR selected mergers with a sample of 7 optically selected mergers chosen from the “Toomre Sequence” (Toomre 1977). We find that the IR selected mergers have a broader range of encounter geometries and slightly less tidal HI, suggesting that the optically selected systems are biased towards prograde-prograde encounters. Spin geometry therefore does not seem to be playing a deciding role in whether or not a merger experiences an IR bright phase.

We find that both IR and optically selected mergers experience a rapid truncation at column densities of order 10^{19} cm^{−2}, similar to what is found in normal disk spirals and believed to be due to ionization by the intergalactic UV field. We also find several striking examples of HI-optical tidal displacements, which we suggest are due to an interaction between an expanding galactic superwind and the outer tidal HI.

Observed Systems

Infrared Selected Mergers

We observed the nearest and most luminous galaxies from the IRAS Bright Galaxy Survey (Soifer et al. 1987).

— Distance < 100 Mpc; 12 >Log(L_{IR}) >11.5 N=5: Arp 299, NGC 1614, NGC 2623, VV 114, IC 883 — Distance < 150 Mpc; Log(L_{IR}) >12 N=3: Arp 220, Mrk 273, Mrk 231 (Mrk 231 undetected)

FIGURE 1 presents the H i column density contoured upon a spectral representation of deep R-band imaging from the UH88′′ (blue=faint, white=bright). Dotted contours indicate H i absorption against a circumnuclear starburst.

FIGURE 2 presents the HI column density contoured upon a spectral representation of H i line-of-sight velocities (red=positive with respect to systemic, blue=negative with respect to systemic). Dotted contours as above.

Optically Selected Mergers

We selected 7 systems from the “Toomre Sequence”, an optically selected sample of relatively isolated galaxies with two well defined tidal tails chosen from the NGC catalog

(Toomre, 1977).

— Five from Hibbard & van Gorkom 1996 Arp 295, NGC 4676, NGC 520, NGC 3921, NGC 7252 — New HI observations of another Toomre object NGC 4038/9 (see figures at end of this poster)

— Final Toomre object taken from the literature NGC 3256 English et al. 1998.

Comparison Between Optically and IR Selected Mergers:

• Gas Content

FIGURE 3 shows a comparison of the cold gas contents of the two samples (IR selected in red, optically selected in blue). This shows:

— The optically selected mergers are just as gas rich as IR selected, when one considers the total cold gas phases (top panel).

— The optically selected have slightly more tidal HI (middle and bottom panels).

• Hubble Types of Progenitor disks

A comparison of the H i distribution and kinematics suggest that both the optically and IR selected samples result primarily from collisions between two gas-rich disks. In particular, 5/7 of each sample appear to have two kinematically distinct gas-rich systems.

Exceptions: NGC 3921 & NGC 4038/9 from the Toomre Sequence appear to include one gas-rich and one gas-poor progenitor. Due to the lack of a clear second kinematic system, the nature of the second progenitor is indeterminate for Arp 220 & Mrk 273 from the IR selected sample.

• Spin Geometries

The H i kinematics conform (at least to first order) to the simple tidal dynamics expected for merging galaxies: tidally expanding streams of gas and stars with smooth runs of velocity, connecting smoothly back to the progenitors or remnants (see FIGURE 2).

These kinematics can be used to determine the spin geometry of the progenitor disks with respect to the orbital geometry. TABLE 1 shows that the IR selected mergers span a broader range of encounter geometries.

TABLE 1. Encounter Spin Geometries.

Optically Selected Mergers  Spin Geometry  IR Selected Mergers
NGC 4038/9, NGC 4676  Prograde — Prograde  NGC 2623, IC 883

NGC 3256, NGC 7252

NGC 520, NGC 3921  Prograde — ???  Mrk 273
Arp 295  Prograde — Retrograde Arp 299, NGC 1614
Retrograde — Retrograde  VV 114
??? — ???  Arp 220

Prograde = spin and orbital angular momentum have the same sign Retrograde = spin and orbital angular momentum have the opposite sign ??? = spin geometry cannot be determined from the data.

Comparison Between Optically and IR Selected Mergers: Conclusions

The above properties can likely be accounted for by suggesting that the Toomre Sequence is biased towards prograde-prograde encounters^{‡}. This is not surprising, given the double-tailed selection criterion imposed by Toomre. Such encounters will raise longer, more gas-rich tidal features. Therefore, there does not seem to be a specific requirement on the encounter spin geometry in order for a merger to result in an IR luminous phase.

However there may still be more subtle interplays between the progenitor properties and the strength of the resulting starburst, e.g. between progenitor gas fraction, spin geometry, and impact parameter. We will test this by developing specific numerical models of these systems (and would welcome any help!).

‡The one retrograde encounter from the optical sample, Arp 295, is actually not in Toomre’s Sequence but was added by Hibbard (1995) because it was modeled in Toomre & Toomre 1972 and included in their earlier list of merger candidates. It is actually a tail-bridge system, not a tail-tail system.

Truncation of HI at Low Column Density:

FIGURE 4 shows the area subtended by H i above a given column density for the systems observed with the best sensitivity. The last points plotted are observed with a signal-to-noise ratio >3.

This plot shows that the H i column density drops quickly below column densities of order 3 × 10^{19} cm^{−2}, similar to what is found in normal disk spirals (e.g. van Gorkom 1993) and believed to be due to ionization by the intergalactic UV field (e.g. Mahoney 1993, Dove & Shull 1993).

Since the tidal gas falls to smaller column density as it is drawn out over a larger area, we fully expect to find tidal H i at much lower column density than in the preencounter disks. Therefore, if the H i in the progenitor disks extended to similar column densities as seen in undisturbed spirals (i.e. ∼ 3 × 10^{19} cm^{−2}) this is strong evidence that some process removes the low column density material. An ambient UV bath seems the most likely explanation.

Observed tidal extent is a lower limit to the true extent of these systems

Unexpected Results

In several cases, extensive H i tails were discovered where there were no known optical counterparts. Some low surface brightness optical counterparts have subsequently been imaged, but some seem devoid of starlight to μ_{R}>27 mag arcsec^{−2}. FIGURE 5 shows the most striking examples: NGC 520, Arp 299, Arp 220. The greyscale shows the H i column density, while the optical light is contoured in blue. These figures shows that not only are there extensive regions of tidal H i without optical counterparts, but the tidal H i and optical features appear to be anti-correlated: the peak H i column densities lie to one side of brighter optical regions.

These three system have three other properties in common:

(1) They each exhibit signs of the “blow-out” phenomena, where a strong galactic wind appears to have blown through the ISM surrounding the circumnuclear starburst and is expanding outward: Hα plumes in NGC 520 (Hibbard & van Gorkom 1996); ROSAT HRI (Heckman et al. 1996) and Hα plumes (Heckman, Armus & Miley 1990) in Arp 220; ROSAT PSPC plume in Arp 299 (Armus et al. in preparation).

(2) In each case, the axis of “blow-out” is within ∼10^{o} of the location of H i displacements or minima. (see overlay for Fig. 5)

(3) Each system shows evidence for a gas-rich rotating disk in the inner regions, which can serve to funnel the starburst energy along the minor axis.

For these reasons, we suspect that the reason for the displacements is primarily due to the interacting between cold tidal H i and a hot expanding wind fluid.

However, there may be more than one process at work (e.g. the end of Arp 299’s tail). Either way, this cautions us that non-gravitational forces are playing an important role in the dynamics of some merging systems.

New VLA Observations of NGC 4038/9: “The Antennae”

Hibbard, van der Hulst & Barnes, in preparation

• VLA D-array, June 1996, VLA C-array, June 1997

FIGURE A presents the H i column density distribution contoured upon a spectral representation of a combined

B, V, R mosaic (from Hibbard 1995, PhD Thesis). The H i map

was made with nearly natural weighting (Robust parameter of 1) and has a resolution of 30′′ (FWHM). NGC 4038/9 appears to be result of a prograde-prograde merger between a gas-rich and a much more gas-poor disk, giving rise to the southern and northern tails, respectively (see also van der

Hulst 1978).

FIGURE B presents the H i contoured on both optical and near-IR CCD images. On the left, the underlying image is again a spectral representation of the combined B, V, R mosaic. To the right, the underlying image is a spectral representation of a near-IR K′ (2.15 μm) image obtained with the UH 88′′ telescope. The H i map was made with nearly uniform weighting (Robust parameter of -1) and has a resolution of 11′′×8′′ (FWHM). This image shows that the H i is actually anti-correlated with the star formation knots

seen so clearly by HST (Whitmore & Schweizer 1995, Miller et

al. 1998: see postcard). The H i is very closely associated with the dustlanes.

FIGURE 1 FIGURE 2 FIGURE 3 FIGURE 4 FIGURE 5 FIGURE 6 FIGURE A FIGURE B